# X-ray Emission Signatures of Neutron Star Mergers

### v1

@@ 3.1 Prompt jet emission
Observationally, there is a correlation between Lorentz factor and isotropic energy of GRBs with $\Gamma\propto E_{\rm iso}^{1/4}$ (Liang et al., 2010; Ghirlanda et al., 2011). Assuming that different $E_{\rm iso}$ is essentially a result of observing a quasi-universal structured jet at different viewing angles (e.g. Zhang & Mészáros, 2002), we adopt a Lorentz factor profile that follows a relation of $\Gamma(\theta)\propto\epsilon_{\gamma}^{1/4}(\theta)$, where variations in $\Gamma$ and $\epsilon_{\gamma}$ arise solely from differences in viewing angle. The Lorentz factor profile then writes
$\Gamma(E_{\gamma,\,\rm iso})=\Gamma_{0}\,(E_{\gamma,\,\rm iso}/10^{52}\,\rm{erg\,s}^{-1})^{1/4}+1,$ (14)
where $\Gamma_{0}$ is the normalization constant and the $+1$ term guarantees $\Gamma>1$. Liang et al. (2010) and Ghirlanda et al. (2011) used observations of Type-II GRBs to find $\Gamma_{0}=182$. Although the corresponding value of $\Gamma_{0}$ for Type-I GRBs remains uncertain, we adopt the same normalization as a phenomenological approximation. This choice is motivated by the lack of robust observational constraints for Type-I GRBs and the expectation that relativistic dynamics among GRBs remain similar.
Consider a structured GRB jet and two observers, one directly on the jet axis $(\theta,\phi)=(0,0)$ and one off the jet axis by an angle $\theta_{v}$ with corresponding coordinates ($\theta_{0},\phi_{0}$). The on-axis observer sees the brightest portion of the jet, while the off-axis observer sees a much dimmer portion. Both observers can detect emission from surrounding regions of the jet, corrected for relativistic (Doppler) beaming effects.
The Doppler factor can generally be written as
$\mathcal{D}\equiv\frac{1}{\Gamma(1-\beta\cos\theta_{v})}.$ (15)
The on- and off-axis observers of a structured GRB jet will each calculate a particular Doppler factor, $\mathcal{D}_{\rm on}$ and $\mathcal{D}_{\rm off}$, respectively. The ratio of Doppler factors is just $\mathcal{R_{\mathcal{D}}}\equiv\mathcal{D_{\rm off}}/\mathcal{D_{\rm on}}$. More generally, an observer with an arbitrary LoS $(\theta_{0},\phi_{0})$, will calculate a unique $\mathcal{R}_{\mathcal{D}}$ for every other region at $(\theta_{i},\phi_{j})$,
$\mathcal{R_{\mathcal{D}}}(\theta_{0},\phi_{0},\theta_{i},\phi_{j})=\frac{\mathcal{D}(\theta_{0},\phi_{0})}{\mathcal{D}(\theta_{i},\phi_{j})}=\frac{1-\beta}{1-\beta\cos\theta_{w}}<1,$ (16)
where $\theta_{w}$ is the angular distance between $(\theta_{0},\phi_{0})$ and $(\theta_{i},\phi_{j})$. $\mathcal{R}_{\mathcal{D}}$ relates observed jet properties between the on- and off-axis observers,
$dt_{\text{off}}$ $=\mathcal{R_{\mathcal{D}}}^{-1}\,dt_{\text{on}}$ (17)
$E_{\text{off}}$ $=\mathcal{R_{\mathcal{D}}}\,E_{\text{on}}$ (18)
$d\Omega_{\text{off}}$ $=\mathcal{R_{\mathcal{D}}}^{-2}\,d\Omega_{\text{on}}.$ (19)
Numerically, we adopt a $500\times 100$ $(\theta,\phi)$ grid across $2\pi$ steradians in spherical coordinates, with finer spacing near the jet axis at $\theta=0$. This choice of resolution ensures an optimal balance between performance and accuracy in the following analyses (see Fig. 3). We define a Gaussian GRB jet with intrinsic $\epsilon_{\gamma}$ (Eq. 13) and $\Gamma$ (Eq. 14) profiles. In the Trapped zone, $\epsilon_{\gamma}(\theta>\theta_{\rm cut})=0$ and $\Gamma(\theta>\theta_{\rm cut})=1$.
Consider again the observer viewing the jet from an arbitrary LoS with corresponding coordinates $(\theta_{0},\phi_{0})$. The effective energy per solid angle that the observer would detect from all emitting regions ($\theta<\theta_{\rm cut}$) is
$\bar{\epsilon}_{\gamma}(\theta_{0},\phi_{0})=k_{0}\frac{\sum_{i,j}\epsilon_{\gamma}(\theta_{i},\phi_{j})\mathcal{R}_{\mathcal{D}}^{3}\Delta\Omega_{i,j}}{\sum_{i,j}\mathcal{R}_{\mathcal{D}}^{2}\Delta\Omega_{i,j}},$ (20)
where $k_{0}$ is the normalization factor, defined such that an on-axis observer detects an isotropic-equivalent energy in gamma-rays, $4\pi\bar{\epsilon}_{\gamma}(0,0)$, equal to a given (observed) $E_{\rm\gamma,\,iso}$. Note that $\bar{\epsilon}_{\gamma}(\theta_{0},\phi_{0})$ is composed of energy per solid angle along the LoS and Doppler-shifted energy per solid angle from surrounding patches—the observed energy per solid angle follows $\bar{\epsilon}_{\gamma}(\theta_{i},\phi_{j})\propto\epsilon(\theta_{i},\phi_{j})\mathcal{R}_{\mathcal{D}}^{3}$. Fig. 4 maps the observed energy per solid angle.
Figure 4: Maps of observed energy per solid angle, $\epsilon\mathcal{R}_{\mathcal{D}}^{3}$, for a $\theta_{c}=5^{\circ}$ Gaussian GRB jet (top) and isotropic wind (bottom) for an observer at $\theta_{v}=0^{\circ}$ (left), $\theta_{v}=20^{\circ}$ (middle), and $\theta_{v}=40^{\circ}$ (right), with the observer’s position marked by the white cross. The gray region indicates the Trapped zone, where $\theta>\theta_{\rm cut}=35^{\circ}$. The top panels show 1d diameter slices through $\theta_{v}$, indicated by the vertical dashed line.
At each grid $(\theta_{i},\phi_{j})$, one can define an on-grid observer-frame spectrum, ${\partial N_{i,j}(E|E_{p})}/{\partial\Omega}=\mathcal{N}_{i,j}(E|E_{p})$ (e.g., a Band function, Eq. 11). The notation $(E|E_{p})$ indicates that energy, $E$, is not directly Doppler-shifted but indirectly scaled through $E_{p}$, which is defined according to the empirical Amati relation, $E_{p}\propto E_{\gamma,\,\rm iso}^{a}$, where $E_{\gamma,\,\rm iso}=4\pi\bar{\epsilon}_{\gamma}(\theta_{w})$. Classically, for Type-II GRBs, $a=0.5$ (Amati et al., 2002). Later, it was found that Type I GRBs also have a similar correlation but with a track above the Type II branch (Zhang et al., 2009). A recent analysis (Minaev & Pozanenko, 2020) gives $a=0.4$ for the short (Type I GRBs), which is adopted in this work. As we will demonstrate later, assigning a Band function spectrum to each grid point is consistent with the Amati relation near $\theta_{v}\sim\theta_{c}$—the regime wherein the relation was originally established—and deviates at larger $\theta_{v}$. while still consistent with GRB 170817A within uncertainties of the Amati relation, which itself is an outlier to the classical Amati relation (see Fig. 7).
@@
[…]
Then, one can model an intrinsic gamma-ray (or X-ray) luminosity light curve per solid angle $\mathcal{L}_{i,j}={\partial L_{i,j}}/{\partial\Omega}$ at each grid (e.g., a FRED profile; Eq. 10) such that its time integral equals the total gamma-ray (or X-ray) energy per solid angle:
[…]
For the observer at $(\theta_{0},\phi_{0})$, the spectra of other patches $(\theta_{i},\phi_{j})$ appear Doppler-shifted,
$\mathcal{N}_{i,j,\rm off}(E)=\mathcal{N}_{i,j}(E|E_{p}(\theta_{w}))*\mathcal{R}_{\mathcal{D}}^{3}(\theta_{\rm obs},\phi_{\rm obs},\theta_{i},\phi_{j}).$ (24)
$\mathcal{N}_{\rm tot}(E)=\frac{\sum_{i,j}\mathcal{N}_{i,j,\,\rm off}(E)\mathcal{R}_{\mathcal{D}}^{3}\Delta\Omega_{i,j}}{\sum_{i,j}\mathcal{R}_{\mathcal{D}}^{2}\Delta\Omega_{i,j}}.\$ (25)
Figure 5: Total observed spectra (solid) compared to reference Band functions (dashed) with $\alpha=-1$ and $\beta=-2.3$ for various $\theta_{v}$. The reference Band functions are normalized such that the peaks are aligned to the observed spectra.
Similarly the duration and luminosity per solid angle of each patch are also Doppler-shifted,
$\Delta t_{i,j,\rm off}$ $=\Delta t/\mathcal{R}_{\mathcal{D}}(\theta_{\rm obs},\phi_{\rm obs},\theta_{i},\phi_{j})$ (26)
${\mathcal{L}_{i,j,\,\rm off}}$ $={\mathcal{L}_{i,j}}\mathcal{R}_{\mathcal{D}}^{2}(\theta_{\rm obs},\phi_{\rm obs},\theta_{i},\phi_{j}).$ (27)
The total observed light curve per solid angle is then
$\mathcal{L}_{\rm tot}=\frac{\sum_{i,j}{\mathcal{L}_{i,j,\,\rm off}}\mathcal{R}_{\mathcal{D}}^{4}\Delta\Omega_{i,j}}{\sum_{i,j}\mathcal{R}_{\mathcal{D}}^{2}\Delta\Omega_{i,j}}.$ (28)
The total observed duration, $\Delta t_{\rm tot}$ is obtained by interpolating the individual durations $\Delta t_{i,j,\,\rm off}$ across the surface. The X-ray peak luminosity at different viewing angles is shown in Fig. 6.
Figure 6: Prompt X-ray emission peak luminosity as a function of viewing angle. Left panel is for BNS mergers with a long-lived magnetar central engine. Middle panel is for BNS mergers without a long-lived magnetar central engine. Right panel is for BH-NS mergers. Total energy is fixed such that $E_{\gamma,\,\rm iso}=10^{51}\,\rm erg\,$ for $\theta_{c}=5^{\circ}$.
Finally, by applying Eq. (21) to the total observed spectrum (Eq. 25), one obtains the total observed energy per solid angle,
[…]
which directly gives the observed isotropic-equivalent energy, $E_{\rm iso}=4\pi\bar{\epsilon}$, in any $E_{1}-E_{2}$ energy band. By the same token, the isotropic-equivalent luminosity is $L_{\rm iso}=4\pi\mathcal{L}_{\rm tot}$.
As the observer moves off-axis from the core of a structured jet, the local emission along the line of sight ($\theta_{v}$) diminishes significantly relative to the contribution from the bright jet core. For viewing angles $\theta_{v}>\theta_{c}$, most of the observed emission originates from an intermediate region between the LoS and jet axis–effectively the “brightest” visible region (see Fig.4). As a result, even when the intrinsic energy distribution is modeled with a Gaussian profile, the observed isotropic-equivalent energy $E_{\rm iso}(\theta_{v})$ remains elevated at large viewing angles $\theta_{v}\gg\theta_{c}$ (see Fig. 3).
In our model we assign each grid an intrinsic spectrum $\mathcal{N}_{i,j}(E\,|\,E_{p})$ with a peak energy $E_{p}$ determined from the local isotropic-equivalent energy $E_{\rm iso}$ via the Amati relation. This setup ensures that emission near the jet axis (small $\theta_{v}$) adheres to the classical Amati relation. However, for observers at large viewing angles ($\theta_{v}\gg\theta_{c}$), the local emission is faint, and the observed spectrum is dominated instead by Doppler-shifted contributions from the brighter jet core. Since the jet core possesses both higher $E_{p}$ and $\bar{\epsilon}$, the integrated spectrum is biased toward these brighter regions. Consequently, the observed spectrum primarily reflects the properties of the brightest contributing region near the jet core even though the total observed $E_{\rm iso}$ is significantly reduced due to Doppler effects. This leads to a substantially higher observed $E_{p}$ than what would be predicted by the classical Amati relation, causing a noticeable deviation as illustrated in Fig. 7. This same deviation was observed in GRB 170817A, where the measured $E_{p}$ was much higher than predicted by the Amati relation (Goldstein et al., 2017; Zhang et al., 2018a).
Figure 7: Observed isotropic-equivalent energy, $E_{\rm iso}$ (blue, left axis) and observed spectral peak energy $E_{p}$ (red, right axis) as a function of viewing angle $\theta_{v}$. The red dashed line bounded by the red shaded region shows the Amati relation derived by Zhang et al. (2009), $E_{p}\propto E_{\rm iso}^{0.34\pm 0.19}$. The red solid line shows the Amati relation derived by Minaev & Pozanenko (2020) for Type I GRBs, $E_{p}\propto E_{\rm iso}^{0.4}$, which is used in this work. Deviations at larger $\theta_{v}$ occur because the observed spectrum is dominated by emission from the jet core, which maintains a higher intrinsic $E_{p}$ despite strong Doppler deboosting of $E_{\rm iso}$. GRB 170817A data obtained from (Zhang et al., 2018b).
3.2 Wind emission
Specifically for BNS mergers with a long-lived magnetar central engine, there is a late emission component that originates from the nearly isotropic magnetar wind due to the spindown of the magnetar. The X-ray luminosity of the wind is paramterized as
$L_{X}(t)=\eta_{X}L_{\rm sd}(t)$ (30)
where $\eta_{X}$ is the efficiency of converting spindown energy to X-ray luminosity. This factor is poorly known and model dependent (Rowlinson et al., 2014; Gao et al., 2016; Xiao & Dai, 2019). In this work, we adopt $\eta_{X}=0.1$. Using typical SMNS parameters (Table 1) in Eq. (4), we estimate the X-ray luminosity to be $L_{X}\sim 10^{48}\rm\,erg\,s^{-1}$. This corresponds quite well to typical X-ray plateau luminosities and provides a natural explanation for such emission (Lü et al., 2015). For a collapse time $t_{c}=t_{\rm sd}=300\,\rm s$, the total wind energy is $(3\times 10^{51}\rm\,erg)\ \eta_{X,\,-1}^{-1}$.
Numerically, we define an isotropic component with $L_{X}=10^{48}\rm\,erg\,s^{-1}$ and $\Gamma_{\rm wind}=50$. Unlike in Section 3.1, we do not define a spectral model for the X-ray wind. Instead we directly model the on-axis X-ray wind light curve using spindown luminosity (Eq. 5) and calculate $\mathcal{L}_{X,\,i,j,\,\rm off}=\mathcal{L}_{X,\,i,j}\mathcal{R}_{\mathcal{D}}$ for a given observer at ($\theta_{\rm obs},\phi_{\rm obs}$). Then, $\mathcal{L}_{X,\,\rm tot}=\sum_{i,j}\mathcal{L}_{X,\,i,j,\rm\,off}\mathcal{R}_{\mathcal{D}}^{3}\Delta\Omega_{i,j}/\sum_{i,j}\mathcal{R}_{\mathcal{D}}^{2}\Delta\Omega_{i,j}$. Note that as long as ($\theta_{\rm obs},\phi_{\rm obs}$) is within the Free zone, the observer is “on axis” of the wind and $L_{X}$ is unchanged. Once the observer is within the Trapped zone ($\theta>\theta_{\rm cut}$), $\mathcal{R}_{\mathcal{D}}\neq 1$ and $L_{X,\rm\,obs}$ will be attenuated.
3.3 Afterglow
For both BNS and NS-BH mergers, the GRB jet interacts with the circumburst medium and decelerates, producing a pair of shocks: a relativistic forward shock propagating into the unshocked medium and a reverse shock propagating back into the unshocked jet (Mészáros & Rees, 1997, 1999; Sari et al., 1998, 1999; Zhang et al., 2003). Both shocks accelerate electrons that radiate photons across a wide range of energies, from radio to gamma-ray and produce afterglow emission. The reverse shock is only relevant early on during and shortly after it passes through the jet, while the forward shock lasts much longer. The broadband emission produced by relativistic shocks from a GRB jet is termed the GRB afterglow emission. We use VegasAfterglow (Wang et al. 2025 in prep) to model X-ray afterglow light curves for a variety of jet geometries and observer’s LoS.
The more isotropic, non-relativistic kilonova ejecta would interact with the ambient medium to produce an ejecta afterglow. The emission is weak and usually buried beneath the bright GRB jet afterglow emission. It may, however, emerge when the jet enters the deep Newtonian phase at a much later time. The late time X-ray observation of GRB 170817A exhibited a flattening of the light curve—whether it is related to the emergence of the kilonova afterglow is inconclusive (Troja et al., 2020; Hajela et al., 2022). We model this component also using VegasAfterglow (Wang et al. 2025 in prep). We adopt isotropic dynamical ejecta with a total kinetic energy of $\sim 10^{51}$ erg and an average initial velocity of $\sim 0.2$c (Villar et al., 2017; Cowperthwaite et al., 2017; Drout et al., 2017). For a magnetar central engine, energy injection into the dynamical ejecta has also been considered. All other parameters are consistent with those used for the jet afterglow, as listed in Table 2.
4 Results
Table 2: Model parameters for jet, magnetar wind and afterglow.
Model $\epsilon(\theta)$ $\Gamma_{\rm 0}$ $E_{\rm\gamma,iso}$ [erg] $\theta_{c}$ [] $\theta_{\rm cut}$ [] $L_{\rm sd,X}$ [erg/s] $t_{\rm sd}$ [s] $E_{\rm k,iso}$ [erg] $n_{\rm ism}$ [cm-3] $p$ $\varepsilon_{e}$ $\varepsilon_{B}$
BNS-I, BNS-II Gaussian 100 $10^{51}$ 5 35 $10^{48}$ 300 $10^{52}$ $0.05$ 2.15 $10^{-2}$ $10^{-3.7}$
BNS-III, BNS-IV Gaussian 100 $10^{51}$ 5 35 $10^{52}$ $0.05$ 2.15 $10^{-2}$ $10^{-3.7}$
BH-NS w/ TD Gaussian 60 $10^{50}$ 15 60 $10^{51}$ $0.05$ 2.15 $10^{-2}$ $10^{-3.7}$
The evolution and fate of an NS binary and the associated EM signatures depend on the particular pair of compact objects, the post-merger remnant, and the observer’s LoS to the system. For prompt and wind X-ray emission, we use the numerical model from Section 3 to produce prompt light curves for a variety of jet geometries and energetics, central engines, and viewing angles. The modeling parameters for prompt, wind and afterglow are summarized in Table 2.
4.1 BNS Mergers
A BNS merger can progress in several ways, based primarily on the remnant mass and angular velocity (upper panel of Fig. 1). In the following subsections, we adopt the following convention for labeling the evolutionary path based on ascending remnant mass:
  1. 1.
    BNS-I—$M_{\rm rem}<M_{\rm TOV}$: The BNS merger will produce a differentially-rotating remnant (i.e., the HMNS “phase”). Once differential rotation becomes uniform, the remnant will spin down indefinitely as a stable NS.
@@ 4.1 BNS Mergers
    BNS-III—$M_{\rm rem}\sim(1.2-1.3)\,M_{\rm TOV}$: The BNS merger will evolve through the HMNS phase and will quickly collapse into a BH once differential rotation is sufficiently damped by magnetic braking.
@@
    BNS-IV—$M_{\rm rem}\gtrsim 1.3\,M_{\rm TOV}$: The BNS merger will promptly form a BH.
In all models of BNS mergers, we include an opaque ejecta beyond $\theta_{\rm cut}=35^{\circ}$, which is consistent with afterglow modeling of GRB 170817A (Troja et al., 2020). A physical interpretation assumes that the nearly isotropic ejecta is pierced by a jet, creating a narrow hole that is further widened by a magnetar wind preferentially directed along the jet axis (Wang et al., 2024) to $\theta_{\rm cut}=35^{\circ}$.
For each evolutionary path, we produce prompt and afterglow X-ray light curves following the method described in Section 3 (Figs 8 - 10). Observational data of GRB 170817A is over-plotted as a reference, serving as a qualitative check on the model behavior. We emphasize that this comparison is not intended as a fit, but rather as a rough benchmark to assess the general consistency of the BNS merger scenarios.
4.1.1 BNS-I (BNS $\rightarrow$ HMNS $\rightarrow$ stable NS)
Physically, one would expect prompt and afterglow X-rays from the GRB jet and extended X-ray emission from the wind of a long-lived remnant. The extended X-ray emission would not cease suddenly due to collapse but taper off with $L_{\rm sd}\propto t^{-2}$ (Eq. 5). The remnant is expected to be able to emit X-ray signals until it falls beneath the pulsar death line, and even then may still survive indefinitely.
@@ 4.1.1 BNS-I (BNS $\rightarrow$ HMNS $\rightarrow$ stable NS)
  • Free zone: Faint prompt emission and rising afterglow from the GRB jet. The extended wind emission becomes observable once the prompt emission sufficiently diminishes—in some cases the prompt emission will never exceed the wind emission and only a plateau will be observed. See middle panel of Fig. 8.
@@
Figure 8: BNS-I model light curves composed of GRB prompt (red), magnetar wind (blue), GRB afterglow (green), and kilonova afterglow (yellow) emissions. Prompt emission is a fast-rise-exponential-decay model (Eq. 10) with peak luminosity corresponding to a Gaussian jet with $\theta_{c}=5^{\circ}$, $E_{\gamma,\,\rm iso}=10^{52}$ erg, and $t=1\,\rm s$. Magnetar wind emission is calculated from spindown luminosity (Eq. 30) and assumed to be isotropic. GRB and kilonova afterglows are calculated using VegasAfterglow. We account for an opaque dynamical ejecta by imposing a cutoff, $\theta_{\rm cut}=35^{\circ}$, beyond which magnetar wind emission along $\theta_{v}$ is blocked until the opacity time, $t_{\rm thin}$ (vertical dotted black line). GBM gamma-ray data was retrieved from  Goldstein et al. (2017) and (Zhang et al., 2018b); Chandra X-ray data was retrieved from Troja et al. (2017, 2019a, 2019b, 2020).
4.1.2 BNS-II (BNS $\rightarrow$ HMNS $\rightarrow$ SMNS $\rightarrow$ BH)
Physically, one would expect prompt and afterglow X-rays from the GRB jet and extended X-ray emission from the wind of a long-lived remnant. The extended X-ray emission would cease suddenly at $t_{\rm sd}$ (Eq. 6), and only the X-ray afterglow would remain.

### v2

@@ IV Results
The evolution and fate of an NS merger and the associated EM signatures depend on the particular progenitor, post-merger remnant, and observer’s LoS to the system. For X-ray prompt and wind emission, we use the numerical model described in Sections III.1 and III.2, and developed in PromptX (Chen et al., 2025). For X-ray afterglow emission from the jet and kilonova, we use VegasAfterglow (Wang et al., 2025).
GRB 170817A is used as a case study in this work. Using VegasAfterglow, we perform an MCMC simulation for $30,000$ steps and a $30\%$ burn-in on a compiled set of GRB 170817A optical (Lyman et al., 2018; Fong et al., 2019; Lamb et al., 2019), radio (Hallinan et al., 2017; Alexander et al., 2017, 2018; Dobie et al., 2018; Mooley et al., 2018b), and X-ray (Troja et al., 2017, 2019a, 2020) afterglow data.
In our MCMC simulations, the initial Lorentz factor, $\Gamma_{0}$ is formally unbounded, which can lead to nonphysical values if left unrestricted. Following common practice in the literature, we adopt a fiducial value of $\Gamma_{0}=200$, consistent with other studies of GRB 170817A (Ghirlanda et al., 2019; Beniamini et al., 2020; Ryan et al., 2024), to ensure realistic jet dynamics and afterglow modeling. The best-fit model parameters are listed in Table 2, with the corresponding light curves and spectra shown in Fig. 8. The MCMC posterior distribution is presented in Fig. 9.
In the following subsections, we explore various progenitors, central engines and viewing angles to identify the most plausible scenario of GRB 170817A.
Figure 9: MCMC posterior distribution of afterglow parameters from a compiled set of GRB 170817A optical (Lyman et al., 2018; Fong et al., 2019; Lamb et al., 2019), radio (Hallinan et al., 2017; Alexander et al., 2017, 2018; Dobie et al., 2018; Mooley et al., 2018b), and X-ray (Troja et al., 2017, 2019a, 2020) afterglow data
IV.1 BNS Mergers
We adopt the following convention to label the BNS evolutionary path in order of increasing remnant mass:
  1. 1.
@@ IV.1 BNS Mergers
[…]
@@
[…]
In all models of BNS mergers, we postulate an opaque ejecta beyond a cutoff angle, $\theta_{\rm cut}$. This parameter, sometimes called $\theta_{w}$ (wing), is poorly constrained by our MCMC simulations of the afterglow. We adopt $\theta_{\rm cut}=20^{\circ}$ based on the ratio $\theta_{c}:\theta_{v}:\theta_{\rm cut}\approx 1:6:7$ consistent with afterglow modeling of GRB 170817A by  Troja et al. (2019a, 2020); Ryan et al. (2024).
Wang et al. (2024) conducted numerical simulations showing that an initially isotropic wind launched within an ejecta shell interacts with the ejecta and becomes collimated toward the poles. For low ejecta masses ($M_{\rm ej}\sim 10^{-3}M_{\odot}$), this interaction can widen the polar cavity within seconds. Due to limitations on the simulated duration, they were unable to determine the final opening angle of the funnel. Here we assume that in the steady-state Free zone (after $\sim 10^{2}\,\rm s$) the opening angle is $\theta_{\rm cut}=20^{\circ}$.
For each evolutionary path, we show the prompt and afterglow X-ray light curves alongside gamma-ray and X-ray observations of GRB 170817A in Figs. 1012, and discuss the implications of the results.
We note that our results are derived solely from EM observations of the prompt and afterglow emissions following a BNS merger. In the case of GW170817/GRB 170817A, constraints on the merger remnant (e.g., maximum mass and collapse time of an SMNS) have been obtained from the ejecta mass and jet-launching delay timescale assuming that only black holes can power short GRBs (Margalit and Metzger, 2017; Gill et al., 2019). In this paper, we keep open the possibility that SMNSs can power GRB jets, and therefore do not consider these additional constraints.
IV.1.1 BNS-I (BNS $\rightarrow$ HMNS $\rightarrow$ stable NS)
Physically, one expects prompt and afterglow X-ray emission from the GRB jet, as well as extended X-ray emission from the wind of a long-lived remnant. The prompt X-ray emission does not cease immediately after the engine shuts off due to the high-latitude effect Ascenzi et al. (2020), while the extended X-ray emission from the wind decays as $L_{\rm sd}\propto t^{-2}$ (Eq. 8). The kilonova afterglow may be detectable in X-rays at much later times.
@@ IV.1.1 BNS-I (BNS $\rightarrow$ HMNS $\rightarrow$ stable NS)
  • Free zone: Faint prompt emission and rising afterglow from the GRB jet. The extended wind emission becomes observable once the prompt emission sufficiently diminishes—in some cases the prompt emission will never exceed the wind emission and only a plateau may be detected. See middle panel of Fig. 10.
@@
Figure 10: BNS-I model light curves composed of GRB prompt (red), magnetar wind (blue), GRB afterglow (green), and kilonova afterglow (yellow) emissions. Prompt emission is calculated for a Gaussian jet with $\theta_{c}=3^{\circ}$ and $E_{\gamma,\,\rm iso}=10^{51}$ erg (see Section III.1). Magnetar wind emission is calculated from spindown luminosity (Eq. 37) and assumed to be isotropic (see Section III.2). GRB and kilonova afterglows are calculated using VegasAfterglow (see Section III.3). We account for an opaque dynamical ejecta by imposing a cutoff, $\theta_{\rm cut}=20^{\circ}$, beyond which magnetar wind emission along $\theta_{v}$ is blocked until the opacity time, $t_{\rm thin}$ (vertical dotted black line). GBM gamma-ray data was retrieved from  Goldstein et al. (2017) and (Zhang et al., 2018a); Chandra X-ray data was retrieved from Troja et al. (2017, 2019b, 2019a, 2020). The gray shaded region denotes early times without X-ray coverage.
The BNS-I scenario can be definitively ruled out in the case of GRB 170817A due to the lack of X-ray emission signatures until the rising jet afterglow at later times (Troja et al., 2017).
IV.1.2 BNS-II (BNS $\rightarrow$ HMNS $\rightarrow$ SMNS $\rightarrow$ BH)
Physically, one would expect prompt and afterglow X-rays from the GRB jet and extended X-ray emission from the wind of a long-lived remnant. The extended X-ray emission would cease suddenly at $t_{\rm sd}$ (Eq. 9), and only the X-ray afterglow would remain. The kilonova afterglow may be detectable in X-rays at much later times.